A newspaper published an article about a study in which researchers subjected laboratory gloves to stress. Among 285 vinyl​ gloves, 67​% leaked viruses. Among 285 latex​ gloves, 13​% leaked viruses. Using the accompanying display of the technology​ results, and using a 0.05 significance​ level, test the claim that vinyl gloves have a greater virus leak rate than latex gloves. Let vinyl gloves be population 1.

What are the null and alternative​ hypotheses?

Identify the test statistic.

Identify the​ P-value.

What is the conclusion for this​ test?
Answer #1

Given Mi-285 Pi = 0.67 na: 285 -0.13 Hypothesis test Hoipo=P2 (Ho: pisP2) also true. Ha: Pi>P2 pooled proportion P = nipitheP-value p-value with test. Statist`C 2:13:16 is p-value = P(Z >13.16) 0.0000

since P value is less than significance level \alpha = 0.05

so we reject null hypothesis Ho

conclusion

we have sufficient evidence to conclude that vinyl gloves have a greater virus leak rate than latex gloves.